OpenStudy (anonymous):

what property is this . –(7 + 8) = -7 + (-8)

OpenStudy (anonymous):

solve for x here

OpenStudy (anonymous):

need to know what property is used

OpenStudy (baldymcgee6):

solve for x?

OpenStudy (baldymcgee6):

It is the distributive property

OpenStudy (anonymous):

thxs do you know this one . 6 + (-3) + 4 = 4 + 6

OpenStudy (anonymous):

That doesn't look like any legitimate property I've ever seen.

OpenStudy (baldymcgee6):

This can't be a property, it is not even a true statement.

OpenStudy (anonymous):

Typo?

OpenStudy (anonymous):

it asks : Identify and define the property shown in the equation. . 6 + (-3) + 4 = 4 + 6

OpenStudy (anonymous):

That is the property of a statement being false.